i cannot find a problem simliar to this one in my book, could you show me how it is done. Thanks so much
4(2x-1)=16

Answer by stanbon(75887)

4(2x-1)=16
Divide both sides by 4 to get:
2x-1 = 4
Add 1 to both sides to get:
2x=5
Divide both sides by 2 to get:
x=5/2
--------
Cheers,
Stan h